OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2)?

OpenStudy (eric_d):

slope= gradient,m

OpenStudy (eric_d):

Slope/Gradient,m = y2-y1/x2-x1

OpenStudy (anonymous):

1 2 -2 -1

OpenStudy (eric_d):

Plug in the given coordinates into that formula

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

i got 2,-5

OpenStudy (eric_d):

No Let me show you....

OpenStudy (eric_d):

(-4, 4) and (-1, -2)

OpenStudy (eric_d):

Slope/Gradient,m = y2-y1/x2-x1

OpenStudy (eric_d):

m= -2-4 / -1-(-4) =-6/3 =-2

OpenStudy (anonymous):

ohh lol thanks

OpenStudy (eric_d):

You're welcome!

OpenStudy (anonymous):

can you help me with a few more ?

OpenStudy (anonymous):

What type of line is the graph of x = -2?

OpenStudy (anonymous):

A.Vertical B. Horizontal C. Rising D. Falling

OpenStudy (anonymous):

@eric_d